OpenStudy (anonymous):

MEDAL FOR BEST ANSWER! What is the 50th term of the sequence 20,9,-2,-13,...?

rvc (rvc):

-519 is ur ans

rvc (rvc):

is it so?robert

OpenStudy (anonymous):

it is an A.P. a=20 c.d. d=9-20=-11 an=a+(n-1)d n=50 calculate t50